# Iwahori Matsumoto presentation for modules of Iwahori fixed functions on symmetric spaces

### v2

@@ 5.1 The length function
Let $T$ be the maximal torus which corresponds to $\mathcal{A}$. Notice that modulo the action of $T$ the number of chambers is finite (in fact equal to #$\#W$). Thus modulo the action of $T$ the number of pairs $(\mathcal{C}_{1},\mathcal{C}_{2})$ with $\mathcal{C}_{1},\mathcal{C}_{2}$ chambers in $\mathcal{A}$ such that $d(\mathcal{C}_{1},\mathcal{C}_{2})=d$, is finite. Let $\mathcal{C}_{1},\mathcal{C}_{2}$ be two chambers in $\mathcal{A}$ such that $l_{\sigma}(\mathcal{C}_{1})=l_{\sigma}(\mathcal{C}_{2})$ and such that $(\mathcal{C}_{1},\sigma(\mathcal{C}_{1}))$ and $(\mathcal{C}_{2},\sigma(\mathcal{C}_{2}))$ are in the same $T$ orbit. There is $t\in T$ such that $t\mathcal{C}_{1}=\mathcal{C}_{2}$ and $t\sigma(\mathcal{C}_{1})=\sigma(\mathcal{C}_{2})$. This implies that the action of $t$ on $\mathcal{A}$ commutes with $\sigma$ and by the previous proposition $t$ may be replaced with an element of $T\cap H$. We obtain $\mathcal{C}_{2}=t\mathcal{C}_{1}$ with $t\in H$ so $\mathcal{C}_{1}$ and $\mathcal{C}_{2}$ are in the same $H$ orbit. ∎
@@
Let $\mathbf{G}$ be a connected reductive group split over a $p$-adic field $F.$ Let $\mathbf{H}=\mathbf{G}^{\sigma}$ for some algebraic involution $\sigma$ and let $X=\mathbf{G}(F)/\mathbf{H}(F)$ be a $p$-adic symmetric space. Then $S(X)^{I}$ is a finitely generated module over ${\cal H}(G,I).$
@@
Consider the collection $\cal{Y}$ of $H$ orbits of chambers that contain a point fixed by $\sigma$. We claim that $\cal{Y}$ is finite and that the characteristic functions of these chambers (together with some $\Omega$ color) generate $S(X)^{I}$ over ${\cal H}(G,I).$
We begin by showing that $\cal{Y}$ is finite. We know that there is a finite number of $H$ orbits on $\sigma$ stable apartments, so it is enough to fix a $\sigma$ stable apartment $\mathcal{A}$ and to show that in $\mathcal{A}$ the number of $H$ orbits of chamber with a $\sigma$ fixed point is finite. If $\mathcal{C}$ is a chamber that has a $\sigma$ fixed point then $\sigma(\mathcal{C})\cap\mathcal{C}\neq\varnothing$. It is clear that the distance between two intersecting chambers is bounded and thus by the previous proposition there is a finite number of $H$ orbits of such chambers.
In order to prove the second part, we again can fix a $\sigma$ stable apartment $\mathcal{A}$, let $\mathcal{C}$ be a chamber in $\mathcal{A}$ without a $\sigma$ fixed point. It is enough to show that there is a minimal gallery between $\mathcal{C}$ and $\sigma(\mathcal{C})$ such that the chamber $\mathcal{C}_{1}$ which appears immediately after $\mathcal{C}_{0}=\mathcal{C}$ satisfies $l_{\sigma}(\mathcal{C}_{1})=l_{\sigma}(\mathcal{C})-2$. If we will show this we will be able to finish using corollary 5.6, by induction on $l_{\sigma}(\mathcal{C})$.
It is enough to show that there is a point $x\in\mathcal{C}$ such that the line connecting $x$ and $\sigma(x)$ intersects some chamber adjacent to $\mathcal{C}$. Assuming the contrary we will obtain that $\frac{x+\sigma(x)}{2}\in\mathcal{C}$ but $\mathcal{C}$ doesn’t contain $\sigma$ fixed points. A contradiction.
@@
The corollary can be deduced by combining the finite multiplicity result (Theorem 5.7) of [Del10] and the finite type results (Theorem B.0.2) of Appendix B of [AGS15]. We believe our proof is more elementary.
@@ 6 Involutions and parahoric group schemes
In order to deal with the equality case $l_{\sigma}(s\times x)=l_{\sigma}(x)$ we will use reduction to a symmetric space of a group with semi simple rank 1 over a finite field. In order to preform this reduction we consider, in this section, the action of $\sigma$ on certain parahoric group schemes.
Assume that $\mathcal{A}$ is a $\sigma$ stable apartment of $\mathcal{B}_{G}$, let $L\subset\mathcal{A}$ be a face of some chamber in $\mathcal{A}$. Denote by $span(L)$ the minimal affine hyper plane that contains $L$ (this is an affine hyper plane and not a linear one). Assume that $span(L)$ is $\sigma$ stable, and let $L^{\prime}=L\cup\sigma(L)$, clearly $L^{\prime}$ is $\sigma$ stable.
Let ${\cal G}_{L}$, ${\cal G}_{\sigma(L^{\prime})}$ and ${\cal G}_{L^{\prime}}$ be the $\mathcal{O}$ group schemes which corresponds to $L$, $\sigma(L)$ and $L^{\prime}$ respectively (see section 2.3).
@@
$\sigma$ induces an involution $\sigma:{\cal G}_{L^{\prime}}\rightarrow{\cal G}_{L^{\prime}}$, $\sigma$ also induces a map ${\cal G}_{L}\rightarrow{\cal G}_{\sigma(L)}$.
@@
We define the map on the open subscheme of ${\cal G}_{L^{\prime}}$ isomorphic to ${\cal U}^{+}_{L^{\prime}}\times\mathbf{T}\times{\cal U}^{-}_{L^{\prime}}$ (see Theorem 2.24). $\mathbf{T}$ is $\sigma$ stable and for any root $\alpha$, ${\cal U}_{\alpha_{L^{\prime}}}$ is mapped to ${\cal U}_{\sigma(\alpha)_{L^{\prime}}}$. $\sigma$ can be extended to ${\cal G}_{L^{\prime}}$ in a unique way by Theorem 0.10 of [Lan95].
@@
We have restriction maps $res^{L^{\prime}}_{L}:{\cal G}_{L^{\prime}}\rightarrow{\cal G}_{L}$ and $res^{L^{\prime}}_{\sigma(L)}:{\cal G}_{L^{\prime}}\rightarrow{\cal G}_{\sigma(L)}$ (see Proposition 2.27).
@@
${{\cal G}_{L^{\prime}}}$${{\cal G}_{L^{\prime}}}$${{\cal G}_{L}}$${{\cal G}_{\sigma(L)}}$$\scriptstyle{\sigma}$$\scriptstyle{res^{L^{\prime}}_{L}}$$\scriptstyle{res^{L^{\prime}}_{\sigma(L)}}$$\scriptstyle{\sigma}$
@@
On the open subscheme isomorphic to ${\cal U}^{+}_{L^{\prime}}\times\mathbf{T}\times{\cal U}^{-}_{L^{\prime}}$ it is clear. Two maps which commute on an open subscheme must commute on the entire scheme. ∎
Let $C_{L}$ be the set of all chambers that contain $L$.
Let $P_{L}$ be the pointwise stabilizer of $L$, it can be identified with the $\mathcal{O}$ points of ${\cal G}_{L}$.
Let $G_{L^{\prime}}=\overline{\cal G}_{L^{\prime}}/R_{u}(\overline{\cal G}_{L^{\prime}})$ be the special fiber of ${\cal G}_{L^{\prime}}$ divided by its unipotent radical. $G_{L^{\prime}}$ is the reductive quotient associated to $L^{\prime}$. $\sigma$ acts on $G_{L^{\prime}}$.
@@
$H\cap P_{L}=H\cap P_{L^{\prime}}$ and each element of this group projects to a $\sigma$ invariant element of $G_{L^{\prime}}$.
@@
Clearly $H\cap P_{L^{\prime}}\subset H\cap P_{L}$ and if $h\in H$ fixes $L$ pointwise it also fixes $\sigma(L)$ pointwise, meaning we have $H\cap P_{L}=H\cap P_{L^{\prime}}$. The second part is also clear. ∎
@@
Any a $\sigma$ invariant element of $G_{L^{\prime}}$ can be lifted to an element of $H\cap P_{L}$.
@@
Let $x$ be a $\sigma$ invariant element of $G_{L^{\prime}}$. By theorem 2.26 There is a unique Levi factor $M\subset\overline{\cal G}_{L^{\prime}}$ which contains $\mathbf{T}(k)$. $\mathbf{T}(k)$ is $\sigma$ stable and the unipotent radical of $\overline{\cal G}_{L^{\prime}}$ is also $\sigma$ stable, therefore $M$ must be $\sigma$ stable. There is $m\in M$ which projects to $x\in G_{L^{\prime}}$. $\sigma(m)$ also projects to $\sigma(x)=x$ and $\sigma(m)\in M$, therefore we must have $m=\sigma(m)$.
$\overline{\cal G}_{L^{\prime}}$ is smooth and $\overline{\cal G}_{L^{\prime}}^{\sigma}$ is also smooth (see for example Proposition 34 in [Edi92]). By Hansel’s lemma we can find a $\sigma$ invariant element of ${\cal G}_{L^{\prime}}(\mathcal{O})$ which projects to $m$. This element is in $H\cap P_{L^{\prime}}$.
Like in proposition 2.29, consider the following map, which we denote by $\overline{res}_{L}$, from $C_{L}$, the set of chambers that contain $L$, to the flag variety of $G_{L}$. Let $\mathcal{C}\in C_{L}$, $\overline{res}_{L}(\mathcal{C})=Im(\overline{res}^{\mathcal{C}}_{L})$ is a Borel of $G_{L}$.
@@
[…]
Here $B(K)$ denotes the flag variety of a reductive group $K$ over the finite field $k$.
@@
${{\cal G}_{\mathcal{C}}}$${{\cal G}_{\sigma(\mathcal{C})}}$${{\cal G}_{L}}$${{\cal G}_{\sigma(L)}}$$\scriptstyle{\sigma}$$\scriptstyle{res^{\mathcal{C}}_{L}}$$\scriptstyle{res^{\sigma(\mathcal{C})}_{\sigma(L)}}$$\scriptstyle{\sigma}$
@@
Consider the map from $B(G_{L^{\prime}})$ to $C_{L}$ obtained by composition of the maps in the left column of the commutative diagram of the previous proposition. This map is an isomorphism by propositions 2.29 and 2.28. Denote the inverse if this map by $b_{L}:C_{L}\rightarrow B(G_{L^{\prime}})$.
@@
The map $b_{L}$ and its inverse interchange $H\cap Ker(\omega)$ orbits and $G^{\sigma}_{L^{\prime}}$ orbits.
@@
From propositions 6.3 and 6.4 we get that these maps interchange $H\cap P_{L}$ orbits and $G^{\sigma}_{L^{\prime}}$ orbits. Thus we only need to prove that two chambers in $C_{L}$ are in the same $H\cap Ker(\omega)$ orbit if and only if they are in the same $H\cap P_{L}$ orbit.
The first direction $H\cap P_{L}\subset H\cap Ker(\omega)$ follows from the fact that any face contains a special point and that $\omega$ is trivial on element that fix some special point.
@@
For the second direction, it is enough to show that if $\mathcal{C},\mathcal{C}^{\prime}$ are chambers that contain $L$ and $h\in H$ satisfies $h\mathcal{C}=\mathcal{C}^{\prime}$ and $\omega(h)=1$ then $h$ fixes $f$ pointwise.
@@
By item 2 of theorem 2.12 there is an apartment $\mathcal{A}^{\prime}$ which contains $\mathcal{C}$ and $\mathcal{C}^{\prime}$. This apartment corresponds to some maximal torus $T^{\prime}$.
@@
By item 5 of theorem 2.13 there is $n\in N_{G}(T^{\prime})$ which acts on $\mathcal{C}^{\prime}$ and on $\mathcal{C}$ like $h$. $\omega(h)=1$ so we must also have $\omega(n)=1$.
@@
The action of $n$ fixes $\mathcal{C}\cap\mathcal{C}^{\prime}$ pointwise so $h$ fixes $L$ pointwise. Therefore $h\in P_{L}$ which finishes the proof. ∎
@@
On the Spherical building of $G_{L^{\prime}}$ we can define a length function $l_{\sigma}$ in the same way we did for the affine building.
@@
For any chamber $\mathcal{C}\in C_{L}$ the number $l_{\sigma}(b_{L}(\mathcal{C}))-l_{\sigma}(\mathcal{C})$ doesn’t depend on $\mathcal{C}$ and only depends on $L$.
@@
It is enough to check this for chambers which intersect at a face of codimension one. This follows from the results of the next section, in particular proposition 7.5 and corollary 7.11. ∎
@@
The $l_{\sigma}$ lengths on the spherical building can be used to describe the Bruhat order of Borel orbits closures in the Zariski topology. Using the previous proposition we can define a Bruhat order on $C_{L}$.

### v3

@@ 5.1 The length function
Let $T$ be the maximal torus which corresponds to $\mathcal{A}$. Notice that modulo the action of $T$ the number of chambers is finite (in fact equal to #W\#W). Thus modulo the action of $T$ the number of pairs of chambers $(\mathcal{C}_{1},\mathcal{C}_{2})$ in $\mathcal{A}$ such that $d(\mathcal{C}_{1},\mathcal{C}_{2})=d$ is finite. Let $\mathcal{C}_{1},\mathcal{C}_{2}$ be two chambers in $\mathcal{A}$ such that $l_{\sigma}(\mathcal{C}_{1})=l_{\sigma}(\mathcal{C}_{2})$ and such that $(\mathcal{C}_{1},\sigma(\mathcal{C}_{1}))$ and $(\mathcal{C}_{2},\sigma(\mathcal{C}_{2}))$ are in the same $T$ orbit. There is $t\in T$ such that $t\mathcal{C}_{1}=\mathcal{C}_{2}$ and $t\sigma(\mathcal{C}_{1})=\sigma(\mathcal{C}_{2})$. This implies that the action of $t$ on $\mathcal{A}$ commutes with $\sigma$ and by the previous proposition $t$ may be replaced with an element of $T\cap H$. We obtain $\mathcal{C}_{2}=t\mathcal{C}_{1}$ with $t\in H$ so $\mathcal{C}_{1}$ and $\mathcal{C}_{2}$ are in the same $H$ orbit. Thus, the number of $H$ orbits of chambers $\mathcal{C}$ in $\mathcal{A}$ with $l_{\sigma}(\mathcal{C})=d$ is bounded by the number of $T$ orbits of pairs of chambers $(\mathcal{C}_{1},\mathcal{C}_{2})$ with $d(\mathcal{C}_{1},\mathcal{C}_{2})=d$. ∎
@@
Let $\mathbf{G}$ be a connected reductive group split over a $p$-adic field $F.$ Let $\mathbf{H}=\mathbf{G}^{\sigma}$ for some algebraic involution $\sigma$ and let $X=\mathbf{H}(F)\backslash\mathbf{G}(F)$ be a $p$-adic symmetric space. Then $S(X)^{I}$ is a finitely generated module over ${\mathcal{H}}(G,I).$
@@
Consider the collection $\mathcal{Y}$ of $H$ orbits of chambers that contain a point fixed by $\sigma$. We claim that $\mathcal{Y}$ is finite and that the characteristic functions of these chambers (together with some $\Omega$ color) generate $S(X)^{I}$ over ${\mathcal{H}}(G,I).$
We first show that $\mathcal{Y}$ is finite. We know that there is a finite number of $H$ orbits of $\sigma$ stable apartments. It is enough to fix a $\sigma$ stable apartment $\mathcal{A}$ and to show that in $\mathcal{A}$ the number of $H$ orbits of chamber with a $\sigma$ fixed point is finite. If $\mathcal{C}$ is a chamber that has a $\sigma$ fixed point then $\sigma(\mathcal{C})\cap\mathcal{C}\neq\varnothing$. It is clear that the distance between two intersecting chambers is bounded, and thus by the previous proposition there is a finite number of $H$ orbits of such chambers.
In order to prove the second part, we again fix a $\sigma$ stable apartment $\mathcal{A}$. Let $\mathcal{C}$ be a chamber in $\mathcal{A}$ without a $\sigma$ fixed point. It is enough to show that there is a minimal gallery between $\mathcal{C}$ and $\sigma(\mathcal{C})$ such that the chamber $\mathcal{C}_{1}$ which appears immediately after $\mathcal{C}_{0}=\mathcal{C}$ satisfies $l_{\sigma}(\mathcal{C}_{1})=l_{\sigma}(\mathcal{C})-2$. If we show this, we can then conclude using Corollary 5.6, by induction on $l_{\sigma}(\mathcal{C})$.
It is enough to show that there is a point $x\in\mathcal{C}$ such that the line connecting $x$ and $\sigma(x)$ intersects some chamber adjacent to $\mathcal{C}$. Assuming the contrary we can find $x$ such that $\frac{x+\sigma(x)}{2}\in\mathcal{C}$. This is a contradiction as $\mathcal{C}$ does not contain $\sigma$ fixed points.
@@
Theorem 5.9 can be deduced by combining the finite multiplicity result (Theorem 5.7) of [13] and the finite type results (Theorem B.0.2) of Appendix B of [1]. We believe that our proof is more elementary.
@@ 6 Involutions and parahoric group schemes
To deal with the equality case, $l_{\sigma}(x\times s)=l_{\sigma}(x)$, we will use reduction to a symmetric space of a group of semisimple rank 1 over a finite field. To perform this reduction we consider, in this section, the action of $\sigma$ on certain parahoric group schemes.
Assume that $\mathcal{A}$ is a $\sigma$ stable apartment of $\mathcal{B}_{G}$. Let $L\subset\mathcal{A}$ be a facet of some chamber in $\mathcal{A}$. Denote by $span(L)$ the minimal affine hyper plane that contains $L$ (this is an affine hyper plane, not a linear one). Assume that $span(L)$ is $\sigma$ stable, and let $L^{\prime}=L\cup\sigma(L)$, clearly $L^{\prime}$ is $\sigma$ stable.
Let ${\mathcal{G}}_{L}$, ${\mathcal{G}}_{\sigma(L)}$ and ${\mathcal{G}}_{L^{\prime}}$ be the $\mathcal{O}$ group schemes which correspond to $L$, $\sigma(L)$ and $L^{\prime}$ respectively (see Section 2.3).
@@
$\sigma$ induces an involution $\sigma:{\mathcal{G}}_{L^{\prime}}\rightarrow{\mathcal{G}}_{L^{\prime}}$, $\sigma$ also induces a map ${\mathcal{G}}_{L}\rightarrow{\mathcal{G}}_{\sigma(L)}$.
@@
We define the map on the open subscheme of ${\mathcal{G}}_{L^{\prime}}$ isomorphic to ${\mathcal{U}}^{+}_{L^{\prime}}\times\mathbf{T}\times{\mathcal{U}}^{-}_{L^{\prime}}$ (see Theorem 2.25). The torus $\mathbf{T}$ is $\sigma$ stable and for any root $\alpha$, ${\mathcal{U}}_{\alpha_{L^{\prime}}}$ is mapped to ${\mathcal{U}}_{\sigma(\alpha)_{L^{\prime}}}$. The map $\sigma$ can be extended to ${\mathcal{G}}_{L^{\prime}}$ in a unique way by Theorem 0.10 of [21].
@@
We have restriction maps $res^{L^{\prime}}_{L}:{\mathcal{G}}_{L^{\prime}}\rightarrow{\mathcal{G}}_{L}$ and $res^{L^{\prime}}_{\sigma(L)}:{\mathcal{G}}_{L^{\prime}}\rightarrow{\mathcal{G}}_{\sigma(L)}$ (see Proposition 2.28).
@@
${\mathcal{G}}_{L^{\prime}}$${\mathcal{G}}_{L^{\prime}}$${\mathcal{G}}_{L}$${\mathcal{G}}_{\sigma(L)}$$\scriptstyle{\sigma}$$\scriptstyle{res^{L^{\prime}}_{L}}$$\scriptstyle{res^{L^{\prime}}_{\sigma(L)}}$$\scriptstyle{\sigma}$
@@
On the open subscheme isomorphic to ${\mathcal{U}}^{+}_{L^{\prime}}\times\mathbf{T}\times{\mathcal{U}}^{-}_{L^{\prime}}$ it is clear. Two maps that commute on an open subscheme must commute on the entire scheme. ∎
Let $C_{L}$ denote the set of all chambers containing $L$.
Let $P_{L}$ be the pointwise stabilizer of $L$. The group $P_{L}$ can be identified with the $\mathcal{O}$ points of ${\mathcal{G}}_{L}$.
Let $G_{L^{\prime}}=\overline{\mathcal{G}}_{L^{\prime}}/R_{u}(\overline{\mathcal{G}}_{L^{\prime}})$ be the special fiber of ${\mathcal{G}}_{L^{\prime}}$ divided by its unipotent radical. The group $G_{L^{\prime}}$ is the reductive quotient associated to $L^{\prime}$. The involution $\sigma$ acts on $G_{L^{\prime}}$.
We denote by $G^{\sigma}_{L^{\prime}}$ the fixed points of $\sigma$ on $G_{L^{\prime}}$.
@@
We have $H\cap P_{L}=H\cap P_{L^{\prime}}$ and each element of this group projects to a $\sigma$ invariant element of $G_{L^{\prime}}$.
@@
Clearly $H\cap P_{L^{\prime}}\subset H\cap P_{L}$ and if $h\in H$ fixes $L$ pointwise it also fixes $\sigma(L)$ pointwise, hence $H\cap P_{L}=H\cap P_{L^{\prime}}$. The second part is also clear. ∎
@@
Any $\sigma$ invariant element of $G_{L^{\prime}}$ can be lifted to an element of $H\cap P_{L}$.
@@
Let $x$ be a $\sigma$ invariant element of $G_{L^{\prime}}$. By Theorem 2.27 There is a unique Levi factor $M\subset\overline{\mathcal{G}}_{L^{\prime}}$ that contains $\mathbf{T}(k)$. The torus $\mathbf{T}(k)$ is $\sigma$ stable and the unipotent radical of $\overline{\mathcal{G}}_{L^{\prime}}$ is also $\sigma$ stable, therefore $M$ must be $\sigma$ stable. There is $m\in M$ which projects to $x\in G_{L^{\prime}}$. The element $\sigma(m)$ also projects to $\sigma(x)=x$ and $\sigma(m)\in M$, therefore we must have $m=\sigma(m)$.
$\overline{\mathcal{G}}_{L^{\prime}}$ is smooth, and $\overline{\mathcal{G}}_{L^{\prime}}^{\sigma}$ is also smooth (see for example Proposition 34 in [14]). By Hensel’s lemma we can find a $\sigma$ invariant element of ${\mathcal{G}}_{L^{\prime}}(\mathcal{O})$ which projects to $m$. This element lies in $H\cap P_{L^{\prime}}$.
Like in Proposition 2.30, we consider the following map, which we denote by $\overline{res}_{L}$. It is a map from $C_{L}$, the set of chambers that contain $L$, to the flag variety of $G_{L}$. Let $\mathcal{C}\in C_{L}$, define $\overline{res}_{L}(\mathcal{C})=Im(\overline{res}^{\mathcal{C}}_{L})$, is a Borel of $G_{L}$.
We denote by $B(K)$ the flag variety of a reductive group $K$ over the finite field $k$.
@@
$B(G_{L^{\prime}})$$B(G_{L^{\prime}})$$B(G_{L})$$B(G_{\sigma(L)})$$C_{L}$$C_{\sigma(L)}$$\scriptstyle{\sigma}$$\scriptstyle{res^{L^{\prime}}_{L}}$$\scriptstyle{res^{L^{\prime}}_{\sigma(L)}}$$\scriptstyle{\sigma}$$\scriptstyle{\overline{res}^{-1}_{L}}$$\scriptstyle{\overline{res}^{-1}_{\sigma(L)}}$$\scriptstyle{\sigma}$
@@
${\mathcal{G}}_{\mathcal{C}}$${\mathcal{G}}_{\sigma(\mathcal{C})}$${\mathcal{G}}_{L}$${\mathcal{G}}_{\sigma(L)}$$\scriptstyle{\sigma}$$\scriptstyle{res^{\mathcal{C}}_{L}}$$\scriptstyle{res^{\sigma(\mathcal{C})}_{\sigma(L)}}$$\scriptstyle{\sigma}$
@@
Consider the map from $B(G_{L^{\prime}})$ to $C_{L}$ given as composition of the maps in the left column of the commutative diagram of the previous proposition. This map is an isomorphism by Propositions 2.30 and 2.29. Denote the inverse of this map by $b_{L}:C_{L}\rightarrow B(G_{L^{\prime}})$.
@@
The map $b_{L}$ sends $H\cap Ker(\omega)$ orbits to $G^{\sigma}_{L^{\prime}}$ orbits. The inverse of $b_{L}$ sends $G^{\sigma}_{L^{\prime}}$ orbits to $H\cap Ker(\omega)$ orbits. The map $b_{L}$ is a bijection from each orbit onto its image.
@@
By Propositions 6.3 and 6.4 we get that these maps interchange $H\cap P_{L}$ orbits and $G^{\sigma}_{L^{\prime}}$ orbits. Thus we only need to prove that two chambers in $C_{L}$ are in the same $H\cap Ker(\omega)$ orbit if and only if they are in the same $H\cap P_{L}$ orbit.
@@
By definition, $P_{L}\subset Ker(\omega)$. We only have to prove that if $\mathcal{C},\mathcal{C}^{\prime}$ are chambers that contain $L$ and $h\in H$ satisfies $h\mathcal{C}=\mathcal{C}^{\prime}$ and $\omega(h)=1$ then $h$ fixes $L$ pointwise.
@@
By item 2 of Theorem 2.12 there is an apartment $\mathcal{A}^{\prime}$ that contains $\mathcal{C}$ and $\mathcal{C}^{\prime}$. This apartment corresponds to some maximal torus $T^{\prime}$.
@@
By item 5 of Theorem 2.13 there is $n\in N_{G}(T^{\prime})$ which acts on $\mathcal{C}^{\prime}$ and on $\mathcal{C}$ like $h$. We have $\omega(h)=1$ so we obtain $\omega(n)=1$.
@@
The action of $n$ fixes $\mathcal{C}\cap\mathcal{C}^{\prime}$ pointwise so $h$ fixes $L$ pointwise. Therefore $h\in P_{L}$ completing the proof.
We know that $b_{L}$ is a bijection from $C_{L}$ to $B(G_{L^{\prime}})$, in particular it is a bijection from each orbit onto its image. ∎
Example 6.7.
We give some examples to illustrate Proposition 6.6.
  1. 1.
    Let $G=SL_{3}(F)$ and let $\sigma:G\rightarrow G$ be given by $\sigma(g)=(g^{t})^{-1}$. We have $H=G^{\sigma}=SO_{3}(F)$, a symmetric subgroup of $G$. Let $T\subset G$ be the torus of diagonal matrices, $T$ is $\sigma$ stable.
    Let $\mathcal{A}$ be the apartment corresponding to $T$. $\sigma$ acts on $\mathcal{A}$ by reflection across the point corresponding to $SL_{3}(\mathcal{O})$, we denote this point by 0. If we take $L=0$ then we have $\sigma(L)=0$ and $L^{\prime}=L$. In this case $G_{L^{\prime}}=SL_{3}(k)$ and $G_{L^{\prime}}^{\sigma}=SO_{3}(k)$. The map $b_{L}:C_{L}\rightarrow B(G_{L^{\prime}})$ is the map that takes an Iwahori subgroup contained in $SL_{3}(\mathcal{O})$ and reduces it modulo the uniformizer $\pi$ to get a Borel subgroup of $SL_{3}(k)$. The subgroup of $H$ consisting of elements that preserve 0 is $SO_{3}(\mathcal{O})$. The map $b_{L}$ sends $SO_{3}(\mathcal{O})$ orbits to $SO_{3}(k)$ orbits.
    Now, let us choose a different $L$. Let $I$ be the Iwahori subgroup consisting of matrices of the form $\begin{pmatrix}\mathcal{O}^{\times}&\mathcal{O}&\mathcal{O}\\ \pi\mathcal{O}&\mathcal{O}^{\times}&\mathcal{O}\\ \pi\mathcal{O}&\pi\mathcal{O}&\mathcal{O}^{\times}\end{pmatrix}$. Let $P$ be the parahoric subgroup consisting of matrices in $SL_{3}(\mathcal{O})$ of the form $\begin{pmatrix}\mathcal{O}&\mathcal{O}&\mathcal{O}\\ \mathcal{O}&\mathcal{O}&\mathcal{O}\\ \pi\mathcal{O}&\pi\mathcal{O}&\mathcal{O}^{\times}\end{pmatrix}$. Let $L$ be the facet fixed by $P$. The facet $\sigma(L)$ is the facet fixed by the parahoric subgroup consisting of matrices in $SL_{3}(\mathcal{O})$ of the form $\begin{pmatrix}\mathcal{O}&\mathcal{O}&\pi\mathcal{O}\\ \mathcal{O}&\mathcal{O}&\pi\mathcal{O}\\ \mathcal{O}&\mathcal{O}&\mathcal{O}^{\times}\end{pmatrix}$. We have $L^{\prime}=L\cup\sigma(L)$ and its pointwise stabilizer is equal to $P\cap\sigma(P)$. The group $G_{L^{\prime}}$ is the group of matrices in $SL_{3}(k)$ of the form $\begin{pmatrix}k&k&0\\ k&k&0\\ 0&0&k^{\times}\end{pmatrix}$ which is equal to $SL_{3}(k)\cap(GL_{2}(k)\times GL_{1}(k))$.
    The map $b_{L}$ is a map between chambers that contain $L$ to the flag variety of $SL_{3}(k)\cap GL_{2}(k)\times GL_{1}(k)$ which is $\mathbb{P}^{1}(k)$. It sends an Iwahori subgroup contained in $P\cap\sigma(P)$ to its reduction modulo $\pi$, a Borel subgroup of $SL_{3}(k)\cap GL_{2}(k)\times GL_{1}(k)$.
    The involution $\sigma$ acts on $SL_{3}(k)\cap GL_{2}(k)\times GL_{1}(k)$ as transpose inverse and its fixed points are $SL_{3}(k)\cap O_{2}(k)\times O_{1}(k)$. The $H$ orbits on $C_{L}$ are mapped to $SL_{3}(k)\cap O_{2}(k)\times O_{1}(k)$ orbits on $\mathbb{P}^{1}(k)$.
  2. 2.
    We give one more example, this time in type C. We take $G=Sp_{4}(F)$ and we use the following block representation of $Sp_{4}$. A block matrix $\begin{pmatrix}A&B\\ C&D\end{pmatrix}\in Sp_{4}(F)$ if and only if $A^{t}C=C^{t}A,B^{t}D=D^{t}B$ and $A^{t}D-B^{t}C=I_{2}$. Here, $I_{2}$ is the identity $2\times 2$ matrix.
    Let $H=GL_{2}(F)$ the fixed points of an inner involution $\sigma$. The involution $\sigma$ is defined as conjugation by $\begin{pmatrix}1&0&0&0\\ 0&1&0&0\\ 0&0&-1&0\\ 0&0&0&-1\end{pmatrix}$.
    An element $g\in H$ is embedded as $\begin{pmatrix}g&0\\ 0&(g^{t})^{-1}\end{pmatrix}\in Sp_{4}(F)$.
    Let $T$ be the torus of diagonal matrices in $H$. We can think about it as a maximal torus of $G$ using the embedding of $H$ in $G$. Let $\mathcal{A}$ be the apartment corresponding to $T$. On this apartment, $\sigma$ acts trivially.
    We have the Iwahori subgroup consisting of integer matrices of the form $\begin{pmatrix}A&B\\ C&D\end{pmatrix}\in Sp_{4}(\mathcal{O})$, with $A,D$ upper triangular mod $\pi$ and $C$ zero mod $\pi$. We denote this Iwahori subgroup by $I$. One of the parahoric subgroups that contains $I$ is the subgroup $P$ of matrices in $Sp_{4}(\mathcal{O})$ of the form $\begin{pmatrix}\mathcal{O}^{\times}&\mathcal{O}&\mathcal{O}&\mathcal{O}\\ \pi\mathcal{O}&\mathcal{O}&\mathcal{O}&\mathcal{O}\\ \pi\mathcal{O}&\mathcal{O}&\mathcal{O}&\mathcal{O}\\ \pi\mathcal{O}&\mathcal{O}&\mathcal{O}&\mathcal{O}\end{pmatrix}$. Let $L$ be the facet fixed by $P$, we have $L=\sigma(L)=L^{\prime}$ as $\sigma$ acts trivially on $\mathcal{A}$. We have $G_{L^{\prime}}=GL_{1}\times SL_{2}$ and $\sigma$ acts on $G_{L^{\prime}}$ as an inner involution of $SL_{2}$. Thus, $G_{L^{\prime}}^{\sigma}=GL_{1}\times GL_{1}$. The map $b_{L}$ gives a bijection between Iwahori subgroups contained in $P$ and Borel subgroups of $GL_{1}\times SL_{2}$. $H$ orbits are mapped to $GL_{1}\times GL_{1}$ orbits.
@@
We can define a length function $l_{\sigma}$ on the spherical building of $G_{L^{\prime}}$ in the same way we did for the affine building $\mathcal{B}_{G}$.
@@
For any chamber $\mathcal{C}\in C_{L}$ the number $l_{\sigma}(b_{L}(\mathcal{C}))-l_{\sigma}(\mathcal{C})$ depends only on $L$ and not on $\mathcal{C}$.
@@
It is enough to check this for chambers which intersect at a facet of codimension one. This follows from results in the next section, in particular from Proposition 7.6 and Corollary 7.12. ∎
@@
The $l_{\sigma}$ lengths on the spherical building can be used to describe the Bruhat order of Borel orbit closures in the Zariski topology. Using the previous proposition we can define a Bruhat order on $C_{L}$.